OpenStudy (thekidwithnoanswers):

Solve each system by substitution 8x + 6y = -24 -4x - y = 4

OpenStudy (thekidwithnoanswers):

Solve each system by elimination not Substitution

OpenStudy (mrs.ambrose614):

Substitute -4x

OpenStudy (thekidwithnoanswers):

4x+2y=-24?

OpenStudy (sunnnystrong):

Step one: Multiply one equation by factor to cancel variable \[8x+6y=-24\] \[(-4x-y=4)*6\] \[-24x-6x=24\] Step two: Add equations to cancel one variable by elimination \[8x+6y=-24\] \[-24x-6x=24\] \[-16x=0\] \[x=0\]

OpenStudy (sunnnystrong):

Next: plug in x into one equation & solve for y -4x - y = 4 -4(0)-y=4 y=-4

OpenStudy (thekidwithnoanswers):

really thats all this was?

OpenStudy (mrs.ambrose614):

That's what I told you

OpenStudy (thekidwithnoanswers):

I...you know what it doesnt matter.

OpenStudy (sunnnystrong):

@thekidwithnoanswers Yep:D